ctlseqs

Helper Library for Control Sequences, version 0.1.0

CismonX

This manual is for ctlseqs, a helper library for control sequences. Copyright \begin{matrix}c \\ ◯ \end{matrix} 2021 CismonX <admin@cismon.net>

Permission is granted to copy, distribute and/or modify this document under the terms of the GNU Free Documentation License, Version 1.3 or any later version published by the Free Software Foundation; with no Invariant Sections, with no Front-Cover Texts, and with no Back-Cover Texts. A copy of the license is included in the section entitled “GNU Free Documentation License”.

Short Contents

1 Overview of ctlseqs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 2 Helper Macros . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 3 Control Sequence Matching . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 4 Control Sequence Reading . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 5 Tips & Hints . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 6 Example Programs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 A API Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 B General Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 C GNU Free Documentation License . . . . . . . . . . . . . . . . . . . . . . . 12

Table of Contents

1 Overview of ctlseqs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

1.1 Contributing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 1.2 Use Scenarios of ctlseqs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

2 Helper Macros . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 3 Control Sequence Matching . . . . . . . . . . . . . . . . . . . . 3

3.1 Matcher Configuration. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

3.1.1 Patterns . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 3.2 Matching String . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

4 Control Sequence Reading . . . . . . . . . . . . . . . . . . . . . . 6 5 Tips & Hints. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 6 Example Programs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 Appendix A API Reference . . . . . . . . . . . . . . . . . . . . . . . 9 Appendix B General Index . . . . . . . . . . . . . . . . . . . . . . 11 Appendix C GNU Free Documentation License . . 12

1 Overview of ctlseqs

The name “ctlseqs” is an abbreviation of “control sequences”, as defined in section 5.4 of ECMA-48.

As the name suggests, this library focuses on handling control sequences. However, it only cares about the bit combinations, while the actual meaning and implementation of a control sequence is up to the user.

The C API provided by ctlseqs is composed of three major parts: The helper macros, the control sequence matcher, and the control sequence reader. Any of them can be used separatedly or combined, after including the header file ctlseqs.h in a source file.

1.1 Contributing

We welcome any form of contribution to ctlseqs (as well as this manual), including bug reports, patches, etc.

   Source code of ctlseqs is hosted on Savannah (https://savannah.nongnu.org/projects/ctlseqs).

You can contribute to ctlseqs using the bug tracker and patch manager, or discuss with the community using the mailing lists.

1.2 Use Scenarios of ctlseqs

Control sequences, as well as other control functions, were once commonly used in computer terminals. Terminals exchange control information with the host regarding colors, font styles, cursor position, etc., using control functions embedded in normal text. Such physical terminals are no longer used today, however, popular ones like DEC VT100 are widely emulated by modern terminal emulators.

The primary purpose of the ctlseqs library is to provide developers with simple and easy-to-use API for handling control functions, when working on terminal emulators and text-based programs.

However, since there is no de facto standard, control functions used in terminals are largely vendor-specific, and terminal emulators like to add their own private controls. That makes ctlseqs not suitable for writing text-based programs which intend to be portable. Instead of raw control codes, developers should stick to ncurses or terminfo.

There are still cases when dealing with raw escape sequences is inevitable, and ctlseqs may come in handy:

• Development of text-based programs which rely heavily on special control sequences, which is not supported by libraries like ncurses.

• Implementing a terminal emulator.

• Experimenting or debugging the features of text-based programs or terminal emulators.

2 Helper Macros

ctlseqs provides C preprocessor macros representing control functions, which expand to C string literals.

The control function can be one of the following three types:

• Elements from the C0 or C1 set.

• Control Sequences.

• Other control functions (such as device control functions).

The name of a helper macro is the control function name with CTLSEQS_as prefix. For a control function other than elements from the C0 or C1 set, the corresponding helper macro is a function-like macro which may or may not take arguments.

Control sequences listed in the helper macros are primarily excerpted from XTerm’s man-

ual (https://invisible-island.net/xterm/ctlseqs/ctlseqs.html), which may differ

across implementations.

As ctlseqs does not currently support 8-bit controls, 2-character 7-bit codes from the C1 set are used instead of their 1-character 8-bit representation. For example, CTLSEQS_CSI expands to "\x1b[".

The following code snippet is an example usage of helper macros:

     printf(CTLSEQS_BEL);
     printf(CTLSEQS_XTVERSION());
     printf(CTLSEQS_CUP("%d", "%d"), 3, 4);

Keep in mind that the standard output stream is line buffered in a terminal. Either fflush(stdout) after printing, or disable output buffering with setvbuf(stdout, NULL,

_IONBF, 0).

3 Control Sequence Matching

Given a character string, checking whether it matches a control sequence is quite trivial, with only the standard C library:

char const *str /* = ... */;
int row, col;
if (0 == strcmp(str, CTLSEQS_XTVERSION())) {
    // ...
} else if (2 == sscanf(str, CTLSEQS_CUP("%d", "%d"), &row, &col)) {
    // ...
} else {
    // ...
}

However, as the number of possible matches grows, this naive implementation becomes less efficient and harder to maintain.

Such problems can be easily solved by using the control sequence matcher provided by ctlseqs.

The struct ctlseqs_matcher * is a pointer to an opaque type which represents an instance of control sequence matcher. Before using, the matcher should be initialized with

ctlseqs_matcher_init. After used, it should be deallocated with ctlseqs_matcher_free.
struct ctlseqs_matcher *matcher = ctlseqs_matcher_init();
// ...
ctlseqs_matcher_free(matcher);

On rare occurences when ctlseqs fail to allocate enough memory, function ctlseqs_ matcher_init may return NULL. However, it is safe to pass null pointers to ctlseqs_

matcher_free.

3.1 Matcher Configuration

Matcher configuration consists of two parts: the number of matching patterns, and the pattern values. Invoke function ctlseqs_matcher_config to configure a matcher.

struct ctlseqs_matcher *matcher /* = ... */;
char const *patterns[] = {
    // ...
};
struct ctlseqs_matcher_options options = {
    .patterns  = patterns,
    .npatterns = sizeof(patterns) / sizeof(char const *),
};
int result = ctlseqs_matcher_config(matcher, &options);
// ...

Each invocation of ctlseqs_matcher_config on the same matcher overwrites the data generated from the last invocation. Upon success, the function returns CTLSEQS_OK. If the function fails to allocate enough memory, returns CTLSEQS_NOMEM.

Caution: If the patterns field in struct ctlseqs_matcher_options is invalid,

function behaviour is undefined. See Section 3.1.1 [Patterns], page 4, for details.

3.1.1 Patterns

The patterns field in struct ctlseqs_matcher_options is an array of NUL-terminated

strings which indicates the desired patterns of control functions for the current matcher.

The following types of control functions are recognizable by the matcher:

  • Control sequences: CSI [param...] [intmd...] final

• C1 functions with command string: (APC|DCS|OSC|PM) [cmdstr] ST

• Single shifts: (SS2|SS3) ch

  • SOS function: SOS [chrstr] ST

According to ECMA-48, CSI parameter bytes are of range 0x30 to 0x3f, intermediate bytes 0x20 to 0x2f, and final byte 0x40 to 0x7e. Command string consists of printable characters and characters of range 0x08 and 0x0e. Character string can be any bit combination which does not represent SOS or ST.

A supported control function, either verbatim or combined with placeholders, can be specified as a valid pattern. The terminating NUL character does not count into the pattern.

A placeholder indicates that when matching a string against the pattern, the value at the placeholder’s location should conform to its rules. A placeholder can only take place in the param, intmd, cmdstr or chrstr fields, and can be one of the following values:

• CTLSEQS_PH_NUM: An unsigned integer.

• CTLSEQS_PH_NUMS: Multiple unsigned integers separated with the semicolon ASCII character (value 0x3b).

• CTLSEQS_PH_STR: A string of printable characters.

• CTLSEQS_PH_CMDSTR: A string containing only printable characters and characters of

    range 0x08 to 0x0d.

• CTLSEQS_PH_CSI_PARAM: A string of CSI parameter bytes.

• CTLSEQS_PH_CSI_INTMD: A string of CSI intermediate bytes.

• CTLSEQS_PH_HEXNUM: A string representing a hexadecimal number.

• CTLSEQS_PH_CHRSTR: A string of any bit combination which does not represent SOS or

    ST.

The following code is a valid example of patterns:

     const char *patterns[] = {
         CTLSEQS_CUP(CTLSEQS_PH_NUM, CTLSEQS_PH_NUM),
         CTLSEQS_XTVERSION(),
         CTLSEQS_DECRQM("1000"),
         // ...
     };

3.2 Matching String

Function ctlseqs_match matches a given character string to a matcher. The function accepts four arguments: the matcher, the string to match, length of the string to match, and a buffer which stores the match result.

Before matching, a buffer which is large enough to store the match result should be allocated. The buffer is an array of union ctlseqs_value, whose definition is shown below:

union ctlseqs_value {
    char const    *str;
    size_t         len;
    unsigned long  num;
};

If the string contains a recognizable control function, or part of a control function which is not yet terminated by the end of the string, the length of control function will be stored at len field of match result buffer at offset 0, and the pointer to the first character of the control funtion at the str field at offset 1.

If ctlseqs_match fails to find any control functions, returns CTLSEQS_NOSEQ. For a partial control function, returns CTLSEQS_PARTIAL. If the matcher is not configured with a matching pattern of the control function, the function returns CTLSEQS_NOMATCH.

If the control function matches a pattern configured in the matcher, returns the offset of the matched pattern, and stores the extracted values to the result buffer according to each of the placeholders, starting from offset 2:

• CTLSEQS_PH_NUM, CTLSEQS_PH_HEXNUM: The integer is stored in field num.

• CTLSEQS_PH_NUMS: The number of integers is stored in field len, followed by that many integers stored in field num.

• CTLSEQS_PH_CSI_PARAM, CTLSEQS_PH_CSI_INTMD, CTLSEQS_PH_STR, CTLSEQS_PH_

CMDSTR, CTLSEQS_PH_CHRSTR: The length of the string is stored in field len, followed by a pointer to the first character of the string stored in field str.

The following code is an example of invoking ctlseqs_match:

union ctlseqs_value     buffer[4];
struct ctlseqs_matcher *matcher /* = ... */;
// ...
char const *str     = "foo" CTLSEQS_CUP("2", "4");
size_t      str_len = sizeof("foo" CTLSEQS_CUP("2", "4")) - 1;
ssize_t     result  = ctlseqs_match(matcher, str, str_len, buffer);
assert(result == 0);
assert(buffer[0].len == 6 && buffer[1].str == str + 3);
assert(buffer[2].num == 2 && buffer[3].num == 4);

Function ctlseqs_match allows NULL value for argument matcher, in which case it behaves like a matcher configured with zero patterns is provided.

Caution: If the given string can match multiple patterns in the matcher, it is unspecified which one of them will be the final match result.

6

4 Control Sequence Reading

7

5 Tips & Hints 6 Example Programs

In the source code repository of ctlseqs, there are a few programs that demonstrates the basic usage of ctlseqs.

See “example/sixdraw.c” for a simple TUI program in which you can draw lines on the terminal window using a mouse.

See “tests/tcsgrep.c” for a program which parses control functions into human-readable text. It was originally written as a tool to run tests for ctlseqs, but can also be considered

as a crude alternative of GNU Teseq (https://www.gnu.org/software/teseq/).

Appendix A API Reference

This section contains a complete list of functions exposed by ctlseqs. See the corresponding man pages for details.

Initialize  matcher  (https://nongnu.org/ctlseqs/man-pages/man3/
ctlseqs_matcher_init.3.html):
struct ctlseqs_matcher *ctlseqs_matcher_init(void);
Configure  matcher  (https://nongnu.org/ctlseqs/man-pages/man3/
ctlseqs_matcher_config.3.html):
int ctlseqs_matcher_config(
    struct ctlseqs_matcher               *matcher,
    struct ctlseqs_matcher_options const *options
);
Match string (https://nongnu.org/ctlseqs/man-pages/man3/ctlseqs_match.3.
html):
ssize_t ctlseqs_match(
    struct ctlseqs_reader const *matcher,
    char                  const *str,
    size_t                       str_len,
    union ctlseqs_value         *result
);
Destroy matcher (https://nongnu.org/ctlseqs/man-pages/man3/
ctlseqs_matcher_free.3.html):
void ctlseqs_matcher_free(
    struct ctlseqs_matcher *matcher
);
Initialize reader (https://nongnu.org/ctlseqs/man-pages/man3/
ctlseqs_reader_init.3.html):
struct ctlseqs_reader *ctlseqs_reader_init(void);
Configure reader (https://nongnu.org/ctlseqs/man-pages/man3/
ctlseqs_reader_config.3.html):
int ctlseqs_reader_config(
    struct ctlseqs_reader               *reader,
    struct ctlseqs_reader_options const *options
);
Read and match (https://nongnu.org/ctlseqs/man-pages/man3/ctlseqs_read.3.
html):
ssize_t ctlseqs_read(
    struct ctlseqs_reader        *reader,
    struct ctlseqs_matcher const *matcher,
    int                           timeout
);

Purge reader (https://nongnu.org/ctlseqs/man-pages/man3/ctlseqs_purge.3.
html):
void ctlseqs_purge(
    struct ctlseqs_reader *reader,
    size_t                 nbytes
);
Destroy reader (https://nongnu.org/ctlseqs/man-pages/man3/
ctlseqs_reader_free.3.html):
void ctlseqs_reader_free(
    struct ctlseqs_reader *reader
);

Appendix B General Index

C

Control functions supported by the matcher . . . . . 4 Control sequence matcher

configuration example. . . . . . . . . . . . . . . . . . . . . . . . . 3 Control sequence matcher

initialization example . . . . . . . . . . . . . . . . . . . . . . . . . 3 Control sequence matcher pattern example . . . . . . 4

H

Helper macro usage example . . . . . . . . . . . . . . . . . . . . 2

L

List of common use cases of ctlseqs . . . . . . . . . . . . . . 1 List of control function types in

ctlseqs helper macros . . . . . . . . . . . . . . . . . . . . . . . . . 2 List of supported placeholders . . . . . . . . . . . . . . . . . . . 4

Appendix C GNU Free Documentation License

Version 1.3, 3 November 2008

Copyright \begin{matrix}c \\ ◯ \end{matrix} 2000, 2001, 2002, 2007, 2008 Free Software Foundation, Inc.

 https://fsf.org/

Everyone is permitted to copy and distribute verbatim copies of this license document, but changing it is not allowed.

0. PREAMBLE The purpose of this License is to make a manual, textbook, or other functional and useful document free in the sense of freedom: to assure everyone the effective freedom to copy and redistribute it, with or without modifying it, either commercially or noncommercially. Secondarily, this License preserves for the author and publisher a way to get credit for their work, while not being considered responsible for modifications made by others. This License is a kind of “copyleft”, which means that derivative works of the document must themselves be free in the same sense. It complements the GNU General Public License, which is a copyleft license designed for free software. We have designed this License in order to use it for manuals for free software, because free software needs free documentation: a free program should come with manuals providing the same freedoms that the software does. But this License is not limited to software manuals; it can be used for any textual work, regardless of subject matter or whether it is published as a printed book. We recommend this License principally for works whose purpose is instruction or reference.

1. APPLICABILITY AND DEFINITIONS This License applies to any manual or other work, in any medium, that contains a notice placed by the copyright holder saying it can be distributed under the terms of this License. Such a notice grants a world-wide, royalty-free license, unlimited in duration, to use that work under the conditions stated herein. The “Document”, below, refers to any such manual or work. Any member of the public is a licensee, and is addressed as “you”. You accept the license if you copy, modify or distribute the work in a way requiring permission under copyright law. A “Modified Version” of the Document means any work containing the Document or a portion of it, either copied verbatim, or with modifications and/or translated into another language. A “Secondary Section” is a named appendix or a front-matter section of the Document that deals exclusively with the relationship of the publishers or authors of the Document to the Document’s overall subject (or to related matters) and contains nothing that could fall directly within that overall subject. (Thus, if the Document is in part a textbook of mathematics, a Secondary Section may not explain any mathematics.) The relationship could be a matter of historical connection with the subject or with related matters, or of legal, commercial, philosophical, ethical or political position regarding them. The “Invariant Sections” are certain Secondary Sections whose titles are designated, as being those of Invariant Sections, in the notice that says that the Document is released under this License. If a section does not fit the above definition of Secondary then it is not allowed to be designated as Invariant. The Document may contain zero Invariant Sections. If the Document does not identify any Invariant Sections then there are none.

The “Cover Texts” are certain short passages of text that are listed, as Front-Cover Texts or Back-Cover Texts, in the notice that says that the Document is released under this License. A Front-Cover Text may be at most 5 words, and a Back-Cover Text may be at most 25 words.

A “Transparent” copy of the Document means a machine-readable copy, represented in a format whose specification is available to the general public, that is suitable for revising the document straightforwardly with generic text editors or (for images composed of pixels) generic paint programs or (for drawings) some widely available drawing editor, and that is suitable for input to text formatters or for automatic translation to a variety of formats suitable for input to text formatters. A copy made in an otherwise Transparent file format whose markup, or absence of markup, has been arranged to thwart or discourage subsequent modification by readers is not Transparent. An image format is not Transparent if used for any substantial amount of text. A copy that is not “Transparent” is called “Opaque”.

Examples of suitable formats for Transparent copies include plain ASCII without markup, Texinfo input format, LaTEX input format, SGML or XML using a publicly available DTD, and standard-conforming simple HTML, PostScript or PDF designed for human modification. Examples of transparent image formats include PNG, XCF and JPG. Opaque formats include proprietary formats that can be read and edited only by proprietary word processors, SGML or XML for which the DTD and/or processing tools are not generally available, and the machine-generated HTML, PostScript or PDF produced by some word processors for output purposes only.

The “Title Page” means, for a printed book, the title page itself, plus such following pages as are needed to hold, legibly, the material this License requires to appear in the title page. For works in formats which do not have any title page as such, “Title Page” means the text near the most prominent appearance of the work’s title, preceding the beginning of the body of the text.

The “publisher” means any person or entity that distributes copies of the Document to the public.

A section “Entitled XYZ” means a named subunit of the Document whose title either is precisely XYZ or contains XYZ in parentheses following text that translates XYZ in another language. (Here XYZ stands for a specific section name mentioned below, such as “Acknowledgements”, “Dedications”, “Endorsements”, or “History”.) To “Preserve the Title” of such a section when you modify the Document means that it remains a section “Entitled XYZ” according to this definition.

The Document may include Warranty Disclaimers next to the notice which states that this License applies to the Document. These Warranty Disclaimers are considered to be included by reference in this License, but only as regards disclaiming warranties: any other implication that these Warranty Disclaimers may have is void and has no effect on the meaning of this License.

2. VERBATIM COPYING You may copy and distribute the Document in any medium, either commercially or noncommercially, provided that this License, the copyright notices, and the license notice saying this License applies to the Document are reproduced in all copies, and that you add no other conditions whatsoever to those of this License. You may not use technical measures to obstruct or control the reading or further copying of the copies you make or distribute. However, you may accept compensation in exchange for copies. If you distribute a large enough number of copies you must also follow the conditions in section 3. You may also lend copies, under the same conditions stated above, and you may publicly display copies.

3. COPYING IN QUANTITY If you publish printed copies (or copies in media that commonly have printed covers) of the Document, numbering more than 100, and the Document’s license notice requires Cover Texts, you must enclose the copies in covers that carry, clearly and legibly, all these Cover Texts: Front-Cover Texts on the front cover, and Back-Cover Texts on the back cover. Both covers must also clearly and legibly identify you as the publisher of these copies. The front cover must present the full title with all words of the title equally prominent and visible. You may add other material on the covers in addition. Copying with changes limited to the covers, as long as they preserve the title of the Document and satisfy these conditions, can be treated as verbatim copying in other respects. If the required texts for either cover are too voluminous to fit legibly, you should put the first ones listed (as many as fit reasonably) on the actual cover, and continue the rest onto adjacent pages. If you publish or distribute Opaque copies of the Document numbering more than 100, you must either include a machine-readable Transparent copy along with each Opaque copy, or state in or with each Opaque copy a computer-network location from which the general network-using public has access to download using public-standard network protocols a complete Transparent copy of the Document, free of added material. If you use the latter option, you must take reasonably prudent steps, when you begin distribution of Opaque copies in quantity, to ensure that this Transparent copy will remain thus accessible at the stated location until at least one year after the last time you distribute an Opaque copy (directly or through your agents or retailers) of that edition to the public. It is requested, but not required, that you contact the authors of the Document well before redistributing any large number of copies, to give them a chance to provide you with an updated version of the Document.

4. MODIFICATIONS You may copy and distribute a Modified Version of the Document under the conditions of sections 2 and 3 above, provided that you release the Modified Version under precisely this License, with the Modified Version filling the role of the Document, thus licensing distribution and modification of the Modified Version to whoever possesses a copy of it. In addition, you must do these things in the Modified Version: A. Use in the Title Page (and on the covers, if any) a title distinct from that of the

Document, and from those of previous versions (which should, if there were any, be listed in the History section of the Document). You may use the same title as a previous version if the original publisher of that version gives permission.

B. List on the Title Page, as authors, one or more persons or entities responsible for authorship of the modifications in the Modified Version, together with at least five of the principal authors of the Document (all of its principal authors, if it has fewer than five), unless they release you from this requirement.

C. State on the Title page the name of the publisher of the Modified Version, as the

publisher. D. Preserve all the copyright notices of the Document. E. Add an appropriate copyright notice for your modifications adjacent to the other

copyright notices. F. Include, immediately after the copyright notices, a license notice giving the public

permission to use the Modified Version under the terms of this License, in the form shown in the Addendum below.

G. Preserve in that license notice the full lists of Invariant Sections and required Cover

Texts given in the Document’s license notice. H. Include an unaltered copy of this License. I. Preserve the section Entitled “History”, Preserve its Title, and add to it an item

stating at least the title, year, new authors, and publisher of the Modified Version as given on the Title Page. If there is no section Entitled “History” in the Document, create one stating the title, year, authors, and publisher of the Document as given on its Title Page, then add an item describing the Modified Version as stated in the previous sentence.

J. Preserve the network location, if any, given in the Document for public access to a Transparent copy of the Document, and likewise the network locations given in the Document for previous versions it was based on. These may be placed in the “History” section. You may omit a network location for a work that was published at least four years before the Document itself, or if the original publisher of the version it refers to gives permission.

K. For any section Entitled “Acknowledgements” or “Dedications”, Preserve the Title of the section, and preserve in the section all the substance and tone of each of the contributor acknowledgements and/or dedications given therein.

L. Preserve all the Invariant Sections of the Document, unaltered in their text and in their titles. Section numbers or the equivalent are not considered part of the section titles.

M. Delete any section Entitled “Endorsements”. Such a section may not be included

in the Modified Version. N. Do not retitle any existing section to be Entitled “Endorsements” or to conflict in

title with any Invariant Section. O. Preserve any Warranty Disclaimers.

If the Modified Version includes new front-matter sections or appendices that qualify as Secondary Sections and contain no material copied from the Document, you may at your option designate some or all of these sections as invariant. To do this, add their titles to the list of Invariant Sections in the Modified Version’s license notice. These titles must be distinct from any other section titles. You may add a section Entitled “Endorsements”, provided it contains nothing but endorsements of your Modified Version by various parties—for example, statements of peer review or that the text has been approved by an organization as the authoritative definition of a standard. You may add a passage of up to five words as a Front-Cover Text, and a passage of up to 25 words as a Back-Cover Text, to the end of the list of Cover Texts in the Modified Version. Only one passage of Front-Cover Text and one of Back-Cover Text may be added by (or through arrangements made by) any one entity. If the Document already includes a cover text for the same cover, previously added by you or by arrangement made by the same entity you are acting on behalf of, you may not add another; but you may replace the old one, on explicit permission from the previous publisher that added the old one. The author(s) and publisher(s) of the Document do not by this License give permission to use their names for publicity for or to assert or imply endorsement of any Modified Version.

5. COMBINING DOCUMENTS You may combine the Document with other documents released under this License, under the terms defined in section 4 above for modified versions, provided that you include in the combination all of the Invariant Sections of all of the original documents, unmodified, and list them all as Invariant Sections of your combined work in its license notice, and that you preserve all their Warranty Disclaimers. The combined work need only contain one copy of this License, and multiple identical Invariant Sections may be replaced with a single copy. If there are multiple Invariant Sections with the same name but different contents, make the title of each such section unique by adding at the end of it, in parentheses, the name of the original author or publisher of that section if known, or else a unique number. Make the same adjustment to the section titles in the list of Invariant Sections in the license notice of the combined work. In the combination, you must combine any sections Entitled “History” in the various original documents, forming one section Entitled “History”; likewise combine any sections Entitled “Acknowledgements”, and any sections Entitled “Dedications”. You must delete all sections Entitled “Endorsements.”

6. COLLECTIONS OF DOCUMENTS You may make a collection consisting of the Document and other documents released under this License, and replace the individual copies of this License in the various documents with a single copy that is included in the collection, provided that you follow the rules of this License for verbatim copying of each of the documents in all other respects. You may extract a single document from such a collection, and distribute it individually under this License, provided you insert a copy of this License into the extracted document, and follow this License in all other respects regarding verbatim copying of that document.

7. AGGREGATION WITH INDEPENDENT WORKS A compilation of the Document or its derivatives with other separate and independent documents or works, in or on a volume of a storage or distribution medium, is called an “aggregate” if the copyright resulting from the compilation is not used to limit the legal rights of the compilation’s users beyond what the individual works permit. When the Document is included in an aggregate, this License does not apply to the other works in the aggregate which are not themselves derivative works of the Document. If the Cover Text requirement of section 3 is applicable to these copies of the Document, then if the Document is less than one half of the entire aggregate, the Document’s Cover Texts may be placed on covers that bracket the Document within the aggregate, or the electronic equivalent of covers if the Document is in electronic form. Otherwise they must appear on printed covers that bracket the whole aggregate.

8. TRANSLATION Translation is considered a kind of modification, so you may distribute translations of the Document under the terms of section 4. Replacing Invariant Sections with translations requires special permission from their copyright holders, but you may include translations of some or all Invariant Sections in addition to the original versions of these Invariant Sections. You may include a translation of this License, and all the license notices in the Document, and any Warranty Disclaimers, provided that you also include the original English version of this License and the original versions of those notices and disclaimers. In case of a disagreement between the translation and the original version of this License or a notice or disclaimer, the original version will prevail. If a section in the Document is Entitled “Acknowledgements”, “Dedications”, or “History”, the requirement (section 4) to Preserve its Title (section 1) will typically require changing the actual title.

9. TERMINATION You may not copy, modify, sublicense, or distribute the Document except as expressly provided under this License. Any attempt otherwise to copy, modify, sublicense, or distribute it is void, and will automatically terminate your rights under this License. However, if you cease all violation of this License, then your license from a particular copyright holder is reinstated (a) provisionally, unless and until the copyright holder explicitly and finally terminates your license, and (b) permanently, if the copyright holder fails to notify you of the violation by some reasonable means prior to 60 days after the cessation. Moreover, your license from a particular copyright holder is reinstated permanently if the copyright holder notifies you of the violation by some reasonable means, this is the first time you have received notice of violation of this License (for any work) from that copyright holder, and you cure the violation prior to 30 days after your receipt of the notice. Termination of your rights under this section does not terminate the licenses of parties who have received copies or rights from you under this License. If your rights have been terminated and not permanently reinstated, receipt of a copy of some or all of the same material does not give you any rights to use it.

10. FUTURE REVISIONS OF THIS LICENSE The Free Software Foundation may publish new, revised versions of the GNU Free Documentation License from time to time. Such new versions will be similar in spirit to the present version, but may differ in detail to address new problems or concerns.

See https://www.gnu.org/licenses/.

Each version of the License is given a distinguishing version number. If the Document specifies that a particular numbered version of this License “or any later version” applies to it, you have the option of following the terms and conditions either of that specified version or of any later version that has been published (not as a draft) by the Free Software Foundation. If the Document does not specify a version number of this License, you may choose any version ever published (not as a draft) by the Free Software Foundation. If the Document specifies that a proxy can decide which future versions of this License can be used, that proxy’s public statement of acceptance of a version permanently authorizes you to choose that version for the Document.

11. RELICENSING “Massive Multiauthor Collaboration Site” (or “MMC Site”) means any World Wide Web server that publishes copyrightable works and also provides prominent facilities for anybody to edit those works. A public wiki that anybody can edit is an example of such a server. A “Massive Multiauthor Collaboration” (or “MMC”) contained in the site means any set of copyrightable works thus published on the MMC site. “CC-BY-SA” means the Creative Commons Attribution-Share Alike 3.0 license published by Creative Commons Corporation, a not-for-profit corporation with a principal place of business in San Francisco, California, as well as future copyleft versions of that license published by that same organization. “Incorporate” means to publish or republish a Document, in whole or in part, as part of another Document. An MMC is “eligible for relicensing” if it is licensed under this License, and if all works that were first published under this License somewhere other than this MMC, and subsequently incorporated in whole or in part into the MMC, (1) had no cover texts or invariant sections, and (2) were thus incorporated prior to November 1, 2008. The operator of an MMC Site may republish an MMC contained in the site under CC-BY-SA on the same site at any time before August 1, 2009, provided the MMC is eligible for relicensing.

ADDENDUM: How to use this License for your documents

To use this License in a document you have written, include a copy of the License in the document and put the following copyright and license notices just after the title page:

        Copyright (C)  year  your name.
        Permission is granted to copy, distribute and/or modify this document
        under the terms of the GNU Free Documentation License, Version 1.3
        or any later version published by the Free Software Foundation;
        with no Invariant Sections, no Front-Cover Texts, and no Back-Cover
        Texts.  A copy of the license is included in the section entitled ``GNU
        Free Documentation License''.

If you have Invariant Sections, Front-Cover Texts and Back-Cover Texts, replace the “with…Texts.” line with this:

          with the Invariant Sections being list their titles, with
          the Front-Cover Texts being list, and with the Back-Cover Texts
          being list.

If you have Invariant Sections without Cover Texts, or some other combination of the three, merge those two alternatives to suit the situation.

If your document contains nontrivial examples of program code, we recommend releasing these examples in parallel under your choice of free software license, such as the GNU General Public License, to permit their use in free software.